OpenStudy (goformit100):

A rectangular park 60 m long and 40 m wide has two concrete crossroads running in the middle of the park and rest of the park has been used as a lawn. If the area of the lawn is 2109 sq. m, then what is the width of the road? A. 2.91 m B. 3 m C. 5.82 m D. None of these

OpenStudy (mayankdevnani):

Area of the park = (60 x 40) m^2 = 2400 m^2. Area of the lawn = 2109 m^2. Area of the crossroads = (2400 - 2109) m^2 = 291 m^2. Let the width of the road be x metres. Then, 60x + 40x - x^2 = 291 x^2 - 100x + 291 = 0 (x - 97)(x - 3) = 0 x = 3.

OpenStudy (mayankdevnani):

@goformit100

OpenStudy (goformit100):

tHANKS

OpenStudy (mayankdevnani):

weclome!!!!!